OpenStudy (anonymous):

#1 x + 3y + 2z = –16 #2 2x – y + 2z = –15 #3 2x – 2y – z = 1 #1 x + 3y + 2z = –16 x + 2z = -3y – 16 x=-3y – 2z – 16 #2 2(-3y – 2z – 16) – y + 2z = –15 -6y – 4z – 32 – y + 2z = -15 X= -7y – 2z = 17 #3 2(-3y – 2z – 16) – 2y – z = 1 -6y – 4z – 32 – 2y – z = 1 -8y – 5z = -31 -7y – 2z = 17 8y + 5z = 31 Y+3z = 48 This is what I have so far. Can anyone help me with?

OpenStudy (anonymous):

Its Solving a System of Equations with Three Variables

sam (.sam.):

You can solve this using vectors, did you learn that before?

OpenStudy (anonymous):

No i haven't

sam (.sam.):

From (1), \[x + 3y + 2z = –16\] \[x =-3y-2z-16\] Sub (1) to (2), \[2x – y + 2z = –15\] \[2(-3y-2z-16) – y + 2z = –15\] \[z=-\frac{7 y}{2}-\frac{17}{2}\] Sub (z) and (x) to (3), \[2(-3y-2(-\frac{7 y}{2}-\frac{17}{2})-16) – 2y – (-\frac{7 y}{2}-\frac{17}{2}) = 1\] Solving it you'll get y=-1